OpenStudy (anonymous):

If f(1-3^x) = x then find f(-8)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Let y=1-3^x

OpenStudy (anonymous):

I know the answer.

OpenStudy (anonymous):

Now, express x interms of y

OpenStudy (anonymous):

y=1-3^x 3^x=1-y x ln3 =ln(1-y) x= ln(1-y) /ln3

OpenStudy (anonymous):

So, f(x)=ln(1-x) /ln3

OpenStudy (anonymous):

Now, can u find f(-8)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

WELCOME